# NN-ResDMD: Learning Koopman Representations for Complex Dynamics with Spectral Residuals

### v1

@@ Abstract
Analyzing long-term behaviors in high-dimensional nonlinear dynamical systems remains a significant challenge. The Koopman operator framework has emerged as a powerful tool to address this issue by providing a globally linear perspective on nonlinear dynamics. However, existing methods for approximating the Koopman operator and its spectral components, particularly in large-scale systems, often lack robust theoretical guarantees. Residual Dynamic Mode Decomposition (ResDMD) introduces a spectral residual measure to assess the convergence of the estimated Koopman spectrum, which helps filter out spurious spectral components. Nevertheless, it depends on pre-computed spectra, thereby inheriting their inaccuracies. To overcome its limitations, we introduce the Neural Network-ResDMD (NN-ResDMD), a method that directly estimates Koopman spectral components by minimizing the spectral residual. By leveraging neural networks, NN-ResDMD automatically identifies the optimal basis functions of the Koopman invariant subspace, eliminating the need for manual selection and improving the reliability of the analysis. Experiments on physical and biological systems demonstrate that NN-ResDMD significantly improves both accuracy and scalability, making it an effective tool for analyzing complex dynamical systems.
@@ 1 Introduction
In the study of complex dynamical systems, a critical challenge lies in accurately extracting and analyzing long-term behavior in high-dimensional nonlinear systems. Various data-driven methods [1, 2, 3, 4, 5, 6] have been developed to address this challenge, with the Koopman operator [7, 8] framework emerging as a powerful tool due to its ability to globally linearize nonlinear systems. Unlike local linearization methods [9, 10], which approximate dynamics near fixed points, the Koopman operator transforms the entire system into a linear form within an infinite-dimensional space, which allows the use of spectral analysis techniques to study complex dynamics.
Despite its promise, practical computational challenges arise from the infinite-dimensional nature of the Koopman operator. Numerical methods such as Extended Dynamic Mode Decomposition (EDMD) [11] have been developed to approximate the Koopman operator using a finite set of observables, making it possible to extract dynamic modes from data. However, EDMD lacks theoretical guarantees of convergence and may fail to capture the full Koopman spectrum accurately, particularly in large-scale, complex systems.
To address these limitations, Residual Dynamic Mode Decomposition (ResDMD) [12] was introduced, offering convergence guarantees through a spectral residual measure that quantifies how well the estimated Koopman spectrum converges to the true spectrum. By assessing convergence, ResDMD eliminates spurious spectral components that do not represent the system’s true dynamics, enhancing the reliability of spectral estimation. However, ResDMD primarily filters precomputed spectra rather than directly approximating Koopman spectra, limiting its ability to independently refine spectral estimates.
In this paper, we propose Neural Network-ResDMD (NN-ResDMD), which overcomes this limitation by providing a method to directly compute Koopman eigenpairs by minimizing the spectral residual. Additionally, NN-ResDMD employs neural networks to automatically select basis functions, eliminating the need for manual intervention, a common challenge in EDMD-based methods. Through experiments on both toy models and real-world high-dimensional systems, we demonstrate that NN-ResDMD significantly improves accuracy and scalability, making it a practical and effective tool for analyzing complex dynamical systems.
@@ 2 Preliminary on Koopman operator
The Koopman operator $\mathcal{K}$ acts on observables $g\in L^{2}(\Omega,\mu)$ as:
$\mathcal{K}g=g\circ F.$
Although $F$ is nonlinear, the Koopman operator $\mathcal{K}$ is linear, enabling spectral analysis of the system in the infinite-dimension function space.
A key aspect of modern Koopman operator theory is Koopman Mode Decomposition (KMD) [6], which represents system dynamics through its spectral components, i.e. the eigenvalues, Koopman modes, and eigenfunctions. The discrete spectrum is particularly important for insights into long-term behavior, such as periodicity and stability. Our analysis emphasizes these spectral components derived from KMD. Specifically, we seek eigenpairs $(\lambda_{i},\phi_{i})$, where $\lambda_{i}$ are eigenvalues and $\phi_{i}$ are the corresponding Koopman eigenfunctions.
One of the most prominent numerical methods to approximate the Koopman operator and its spectral components is the Extended Dynamic Mode Decomposition (EDMD) method, introduced by [11]. In EDMD, a set of observables (dictionary or basis functions) $\mathbf{\Psi}=[\psi_{1},\dots,\psi_{N_{K}}]$ is selected, and the span of these observables defines the subspace $V_{N_{K}}\coloneqq\text{span}\{\psi_{i}\}_{i=1}^{N_{K}}$. Snapshots of the system’s state are then collected, and the method constructs a finite-dimensional approximation of the Koopman operator by solving a least-squares problem that relates the snapshots of observables. This enables the computation of eigenvalues, eigenfunctions, and Koopman modes. Note that while common choices of dictionary functions are polynomials, Fourier basis, RBF functions, etc., the optimal choice of basis functions is usually unknown a priori and depends heavily on the specific dynamical system.
Given independent and identically distributed data snapshots $\{(x_{i},y_{i})\}_{i=1}^{m}$ with $y_{i}=F(x_{i})$, two matrices $\Psi_{X}$ and $\Psi_{Y}$ are formed by evaluating the dictionary on the data snapshots:
$\Psi_{X}=\begin{bmatrix}\psi_{1}(x_{1})&\dots&\psi_{N_{K}}(x_{1})\\ \vdots&\ddots&\vdots\\ \psi_{1}(x_{m})&\dots&\psi_{N_{K}}(x_{m})\end{bmatrix},\quad\Psi_{Y}=\begin{bmatrix}\psi_{1}(y_{1})&\dots&\psi_{N_{K}}(y_{1})\\ \vdots&\ddots&\vdots\\ \psi_{1}(y_{m})&\dots&\psi_{N_{K}}(y_{m})\end{bmatrix}.$
EDMD computes the Koopman matrix approximation as $K=\Psi_{X}^{\dagger}\Psi_{Y}$, where $\Psi_{X}^{\dagger}$ is the pseudo-inverse of $\Psi_{X}$. The eigenvalues of $K$ provide approximations of the Koopman operator’s spectrum, and the Koopman eigenfunctions $\phi_{i}$ are approximated as $\phi_{i}=\mathbf{\Psi}\mathbf{v}_{i}$, where $\mathbf{v}_{i}\in\mathbb{C}^{N_{K}}$ is the $i$-th eigenvector of $K$.
@@ 3 Koopman Operator Learning
While EDMD approximates the Koopman operator, it suffers from spectral pollution as increasing dictionary sizes introduce spurious eigenvalues, obscuring the system’s true dynamics. Residual Dynamic Mode Decomposition (ResDMD) [12] addresses this by filtering out spurious eigenvalues using spectral residuals but relies on precomputed eigenpairs, inheriting inaccuracies from methods like EDMD.
To overcome these limitations, we propose Neural Network-ResDMD (NN-ResDMD), which directly approximates the Koopman operator and its spectrum by minimizing a ResDMD-specific loss function. Using a Feedforward Neural Network (FNN), NN-ResDMD also optimizes dictionary functions for the Koopman invariant subspace, eliminating the need for manual basis selection.
@@ 3.1 ResDMD Review
Now, suppose we have obtained an eigenpair $(\lambda,\phi)$ of $\mathcal{K}$ from EDMD or other methods [13, 14, 15, 16, 17, 18] where $\lambda\in\mathbb{C}$ and the eigenfunction $\phi$ is expanded in terms of dictionary functions, i.e., $\phi=\mathbf{\Psi}\mathbf{v}=\sum_{i=1}^{N_{K}}\psi_{i}v_{i}\in V_{N_{K}}$ for some $\mathbf{v}\in\mathbb{C}^{N_{K}}$, where $v_{i}$ represents weights of the span. Without loss of generality, we consider $\phi$ has been normalized, i.e., $\|\phi\|_{2}=1$. The accuracy of this eigenpair approximation in the ResDMD framework can be measured by computing its squared relative residual using the dictionary in the following way:
@@
$=\sum_{i,j=1}^{N_{K}}\bar{v}_{i}\left[\langle\mathcal{K}\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}-\lambda\langle\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}-\bar{\lambda}\langle\mathcal{K}\psi_{i},\psi_{j}\rangle_{\mu}+|\lambda|^{2}\langle\psi_{i},\psi_{j}\rangle_{\mu}\right]v_{j},$ (3.1)
where $\bar{v}_{i},\bar{\lambda}$ denote the complex conjugate of $v_{i},\lambda$.
This squared relative residual in (A.1) is the theoretical value that measures the distance between $\phi$ and the eigenspace associated with $\lambda$, especially under the assumption that $\lambda$ is in the discrete spectrum of $\mathcal{K}$. To approximate this residual in practice, we apply the Galerkin approximation [19], which states that as the number of data points $m$ increases, the following limits hold:
@@
$\lim_{m\to\infty}\frac{1}{m}\left[\Psi_{Y}^{*}\Psi_{Y}\right]_{ij}$ $=\langle\mathcal{K}\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}=\langle\psi_{i},\mathcal{K}^{*}\mathcal{K}\psi_{j}\rangle_{\mu},$
where * denotes complex conjugate. Using this approximation, the squared relative residual from (A.1) is approximated as follows (see A.1 for more details):
@@
$\widehat{\textit{res}}(\lambda,\phi)^{2}\coloneqq\frac{1}{m}\mathbf{v}^{*}\left[\Psi_{Y}^{*}\Psi_{Y}-\lambda(\Psi_{X}^{*}\Psi_{Y})^{*}-\bar{\lambda}\Psi_{X}^{*}\Psi_{Y}+|\lambda|^{2}\Psi_{X}^{*}\Psi_{X}\right]\mathbf{v}.$ (3.3)
where (3.3), denoted as $\widehat{\textit{res}}(\lambda,\phi)^{2}$, represents the approximation of the theoretical value in (A.1).
By definition in (A.1), the residual quantifies the deviation from the spectral property, measuring how far the estimated eigenpair is from the true spectrum. In practice, (3.3) is calculated for all precomputed eigenpairs, retaining those with residuals below a threshold. However, while residuals help filter and select valid eigenpairs, they do not improve the accuracy of eigenpair estimation.
3.2 Neural Network-ResDMD
General framework In this section, we present the Neural Network-ResDMD (NN-ResDMD) framework, designed to compute the eigenpairs of the Koopman operator directly using ResDMD-based spectral residuals, as illustrated in Figure 1. The method first determines the optimal dictionary functions by minimizing the total residual $J\coloneqq\sum_{i=1}^{N_{K}}\widehat{\textit{res}}(\lambda_{i},\phi_{i})^{2}$, over all computed eigenpairs $\{(\lambda_{i},\phi_{i})\}_{i=1}^{N_{K}}$. The spectral residual directly impacts the finite-dimensional projection of the Koopman operator and our method minimizes this residual to ensure the learned basis functions adequately capture the Koopman dynamics. This approach allows the construction of the Koopman operator matrix $\tilde{K}$ without relying on external methods or post-processing. Equation (3.5) enables NN-ResDMD to compute eigenpairs directly, improving accuracy compared to ResDMD, which relies on filtering precomputed results from other methods.
In this framework, neural networks parameterize the dictionary functions $\mathbf{\Psi}(x;\theta)$, where $\theta$ represents the network parameters. By minimizing the spectral residual $J$, this approach directly optimizes the dictionary functions towards better approximation of Koopman spectral components, which ensures the learned operator captures the underlying spectral properties of the dynamical system. This is fundamentally different from traditional methods like EDMD, which focus on minimizing prediction errors in the observable space without explicitly considering spectral accuracy. The neural network architecture serves as a flexible function approximator, that allows the framework to adaptively learn the optimal dictionary that minimizes spectral residuals, thereby producing more accurate and reliable Koopman spectral decompositions. This spectral-oriented optimization improves the accuracy of eigenvalues approximations and enhances the quality of the computed eigenfunctions, which leads to better characterization of the system’s dynamic behavior.
Figure 1: (Left) The classical ResDMD and (Right) the Neural Networks based ResDMD methods
From Residual to NN This section explains how neural networks are integrated into the ResDMD framework. In ResDMD, the squared relative residual approximation (3.3) measures how well a computed eigenpair fits the dataset. If the Koopman matrix $K$ is well-approximated by the projected Koopman operator $\mathcal{K}_{N_{K}}$, the total residual $J$ should approach zero as more data is provided. Thus, $J$ can be used as a loss function, and the optimal Koopman matrix $\tilde{K}$ is obtained by minimizing:
$J=\sum_{i=1}^{N_{K}}\widehat{\textit{res}}(\lambda_{i},\phi_{i})^{2}.$ (3.4)
which is equivalent to minimization the following (See A.2 for more details):
$J=\frac{1}{m}\|(\Psi_{Y}-\Psi_{X}K)V\|_{F}^{2}$ (3.5)

### v3

# ResKoopNet: Learning Koopman Representations for Complex Dynamics with Spectral Residuals

@@ Abstract
Analyzing long-term behaviors in high-dimensional nonlinear dynamical systems remains challenging, with the Koopman operator framework providing a powerful global linearization approach, though existing methods for approximating its spectral components often suffer from theoretical limitations and reliance on predefined dictionaries. While Residual Dynamic Mode Decomposition (ResDMD) introduced the spectral residual to assess the accuracy of Koopman operator approximation, it only filters precomputed spectra, which prevents it from fully discovering the Koopman operator’s complete spectral information (a limitation sometimes referred to as the ‘spectral inclusion’ problem). We introduce ResKoopNet (Residual-based Koopman-learning Network), a novel method that addresses this limitation by explicitly minimizing the spectral residual to compute Koopman eigenpairs, which can identify a more precise and complete spectrum of the Koopman operator. This approach provides theoretical guarantees while maintaining computational adaptability through a neural network implementation. Experiments on physical and biological systems demonstrate ResKoopNet’s superior accuracy in spectral approximation compared to existing methods, particularly for systems with continuous spectra and high dimensional, which makes it as an effective tool for analyzing complex dynamical systems.
@@ 1 Introduction
In the analysis of complex dynamical systems, a fundamental challenge is accurately capturing and understanding the long-term behavior of highly nonlinear systems, particularly those with high dimensionality. Various data-driven methods [6, 32, 43, 34, 38, 22, 27, 46, 16] have been developed to address this challenge. Among these approaches, the Koopman operator framework [18, 19] has emerged as a powerful tool due to its ability to globally linearize nonlinear systems. Unlike local linearization methods [13, 12], which only approximate dynamics near fixed points, the Koopman operator transforms the entire system into a linear form within an infinite-dimensional function space, enabling the application of spectral analysis techniques to study complex dynamics.
Despite its promise, practical computational challenges arise from the infinite-dimensional nature of the Koopman operator. Numerical methods such as Extended Dynamic Mode Decomposition (EDMD) have been developed to approximate the Koopman operator using a finite set of observables, which makes it possible to extract dynamic modes from data. Important convergence results for EDMD were established by Korda and Mezić [20], which proved that as the number of samples increases, EDMD converges to the $L^{2}$-orthogonal projection of the Koopman operator on a finite-dimensional subspace. They further showed convergence in the strong operator topology as the dimension of the subspace increases, and that accumulation points of the spectra of EDMD correspond to eigenvalues of the Koopman operator with associated eigenfunctions converging weakly, provided the weak limit is nonzero. However, despite these theoretical advances, EDMD faces notable limitation in spectral approximation. It suffers from spectral pollution, where discretization introduces spurious eigenvalues unrelated to the true operator. Moreover, EDMD struggles to accurately capture continuous spectra, which are crucial for characterizing chaotic and complex systems. As a result, EDMD can miss critical information about the underlying dynamics, particularly in systems exhibiting rich, complex behavior.
To address these limitation, Residual Dynamic Mode Decomposition (ResDMD) [8] was introduced to offer convergence guarantees through a spectral residual that measures how well the estimated Koopman spectra are. By assessing convergence of spectral residual, ResDMD eliminates spurious spectral components that do not represent the system’s true dynamics, which enhances the reliability of spectral estimation. ResDMD also introduces the concept of pseudospectrum, regions where the spectral content is significant even without discrete eigenvalues, which is particularly valuable for analyzing systems with continuous spectra. However, ResDMD only filters precomputed polluted spectra, which limits its ability to independently refine spectral estimates.
In this paper, we propose Residual-based Koopman-learning Network (ResKoopNet), which overcomes these limitations by minimizing spectral residuals over eigenpairs while exploring the optimal dictionary space. ResKoopNet employs neural networks to optimize dictionary functions for the Koopman invariant subspace, which enables both discrete eigenvalue computation and pseudospectrum analysis. Through experiments on both simulation models and high-dimensional real-world systems, we demonstrate its superior accuracy and scalability over other methods.
@@ 2 Preliminary on Koopman Operator
The Koopman operator $\mathcal{K}$ acts on observable $g\in\mathcal{F}$ as:
$\mathcal{K}g=g\circ F,$
[…]
$\quad\quad g(x_{n})=[\mathcal{K}^{n}g](x_{0})$
[…]
where $\varphi_{\lambda}$ are the eigenfunctions of $\mathcal{K}$, $c_{\lambda}$ are expansion coefficients, and $\phi_{\theta,g}$ represents the contribution from the continuous spectrum. $[-\pi,\pi]_{per}$ denotes a periodic interval. The discrete spectrum is particularly important for insights into long-term behavior, such as periodicity and stability, while the continuous spectrum characterizes complex behaviors such as mixing, chaos, and transport phenomena [1, 47, 7]. Our analysis emphasizes finding these spectral components including both discrete and continuous part from KMD.
Extended Dynamic Mode Decomposition (EDMD) [44] is a widely used data-driven method for approximating the Koopman operator and its spectral components in KMD. It selects a set of observables $\{\psi_{i}\}_{i=1}^{N_{K}}$ to form a subspace $V_{N_{K}}=\text{span}\{\psi_{i}\}_{i=1}^{N_{K}}$. The method constructs a finite-dimensional approximation of the Koopman operator by solving a least-square problem based on system snapshots and selected dictionary, allowing the computation of eigenvalues, eigenfunctions, and Koopman modes. While common dictionaries include polynomials, Fourier basis, and RBF functions, selecting an optimal dictionary remains system-dependent and challenging. Given data snapshots $\{(x_{i},y_{i})\}_{i=1}^{m}$ with $y_{i}=F(x_{i})$ and $\{x_{i}\}_{i=1}^{m}$ i.i.d. by $\mu$, two data matrices $\Psi_{X}$ and $\Psi_{Y}$ are formed by evaluating the dictionary on these data points:
$\Psi_{X}\coloneqq\begin{bmatrix}\psi_{1}(x_{1})&\dots&\psi_{N_{K}}(x_{1})\\ \vdots&\ddots&\vdots\\ \psi_{1}(x_{m})&\dots&\psi_{N_{K}}(x_{m})\end{bmatrix},\quad\Psi_{Y}\coloneqq\begin{bmatrix}\psi_{1}(y_{1})&\dots&\psi_{N_{K}}(y_{1})\\ \vdots&\ddots&\vdots\\ \psi_{1}(y_{m})&\dots&\psi_{N_{K}}(y_{m})\end{bmatrix}.$
EDMD computes the approximated Koopman matrix representation by $K=\Psi_{X}^{\dagger}\Psi_{Y}$, where $\Psi_{X}^{\dagger}$ is the pseudoinverse of $\Psi_{X}$. The eigenvalues of $K$ provide approximation of the Koopman operator’s discrete spectrum, and the Koopman eigenfunctions $\phi_{i}$ are approximated as $\phi_{i}=\mathbf{\Psi}\mathbf{v}_{i}$, where $\mathbf{v}_{i}\in\mathbb{C}^{N_{K}}$ is the $i$-th eigenvector of $K$ and $\mathbf{\Psi}\coloneqq[\psi_{1},\dots,\psi_{N_{K}}]$ is a (row) vector of observables from the dictionary.
@@ 3 Koopman Operator Learning
EDMD directly approximates the Koopman operator and its spectral components from data but suffers from spectral pollution, where discretization of the infinite-dimensional operator introduces spurious eigenvalues that have no relation to the true spectrum. Residual Dynamic Mode Decomposition (ResDMD) [8] improves spectral accuracy by filtering out these polluted eigenvalues via spectral residual measurement. However, it only filters precomputed spectra and cannot fully discover the Koopman operator’s complete spectral information, a limitation known as the ‘spectral inclusion’ problem where parts of the true spectrum may be missed entirely.
To address these issues, we propose Residual-based Koopman-learning Network (ResKoopNet), which minimizes spectral residual over eigenpairs while exploring the optimal dictionary space. This approach identifies a more precise and complete spectrum of the Koopman operator while ensuring all learned dictionary functions satisfy small spectral residual criteria. We note that while we employ a simple neural network as our optimizer in this work, the specific choice of optimization tool can be adapted based on the data structure and properties of the dynamical system being studied.
@@ 3.1 ResDMD Review
Now, suppose we have an estimated eigenvalue-eigenfunction pair $(\lambda,\phi)$ of $\mathcal{K}$ where $\lambda\in\mathbb{C}$ and $\phi=\mathbf{\Psi}\mathbf{v}=\sum_{i=1}^{N_{K}}\psi_{i}v_{i}\in V_{N_{K}}$ is expressed in the dictionary functions. Assuming $\phi$ is normalized, i.e., $\|\phi\|_{\mu}=1$, the spectral residual that measures the accuracy of this eigenpair is defined as:
@@
where $\bar{v}_{i},\bar{\lambda}$ denote the complex conjugates of $v_{i},\lambda$.
This spectral residual in Eq. (3.1) quantifies how far the estimated eigenpair deviates from the true spectrum. In practice, we can approximate this residual using data samples through Galerkin methods [5, 8]; more specifically, the following convergence results hold as the number of data points $m\to\infty$:
@@
$\lim_{m\to\infty}\frac{1}{m}\left[\Psi_{Y}^{*}\Psi_{Y}\right]_{ij}$ $=\langle\psi_{i},\mathcal{K}^{*}\mathcal{K}\psi_{j}\rangle_{\mu},$
where * denotes conjugate transpose of matrix or adjoint operator. Using these relationships, the spectral residual in Eq. (3.1) can be approximated as (see A.2 for more details):
@@
$\begin{split}\widehat{\textit{res}}(\lambda,\phi)^{2}\coloneqq\frac{1}{m}\mathbf{v}^{*}[\Psi_{Y}^{*}\Psi_{Y}-\lambda(\Psi_{X}^{*}\Psi_{Y})^{*}\\ -\bar{\lambda}\Psi_{X}^{*}\Psi_{Y}+|\lambda|^{2}\Psi_{X}^{*}\Psi_{X}]\mathbf{v}.\end{split}$ (3.3)
Notice that Eq. (3.3) is calculated only for precomputed eigenpairs. However, the spectral residual serving as a threshold only filters polluted eigenpairs: it does not explore additional accurate eigenpairs or enhance the approximation. For an extreme case, the spectrum of the approximated Koopman operator could be trivial [8, Example 3.1]. This is precisely the spectral inclusion problem we mentioned earlier, where the computed spectrum is constrained to a subset of the true spectrum, but could potentially omit critical spectral components. Addressing this fundamental limitation is a core objective of our work.
Continuous Spectra and Pseudospectrum The concept of pseudospectrum is particularly useful when studying the continuous spectrum of the Koopman operator, as it captures regions where the resolvent norm (i.e., $\|(\mathcal{K}-\lambda I)^{-1}\|$) is large, even in the absence of discrete eigenvalues. In order to compute the pseudospectrum, a set of grid points as candidate spectrum values $\{z_{1},\ldots z_{n_{z}}\}$ is scanned in the complex plane using the spectral residual; specifically, for each grid point $z_{j}\in\mathbb{C}$, we compute the following SVD problem
$\tau_{j}=\min_{\mathbf{v}_{i}\in\mathbb{C}^{N_{k}}}\widehat{\textit{res}}(z_{j},\mathbf{\Psi}\mathbf{v}_{i}),$ (3.4)
which uses the dictionary $\mathbf{\Psi}$ computed from our proposed ResKoopNet method. The approximated $\epsilon$-pseudospectrum containing the continuous spectra is then given by $\{z_{j}:\tau_{j}<\varepsilon\}$. More details on proof of convergence can be found in Colbrook and Townsend [8, Appendix B].
3.2 Residual-based Koopman-learning Network (ResKoopNet)
General framework In this section, we present the Residual-based Koopman-learning Network (ResKoopNet), which minimizes spectral residual to address the spectral inclusion problem and yield a more accurate and complete Koopman spectrum. The overall pipeline is illustrated in Figure 1. ResKoopNet optimizes the dictionary functions by minimizing the total spectral residual,
$J\coloneqq\sum_{i=1}^{N_{K}}\widehat{\textit{res}}(\lambda_{i},\phi_{i})^{2},$
over all computed eigenpairs $\{(\lambda_{i},\phi_{i})\}_{i=1}^{N_{K}}$. This loss directly influences the quality of the finite-dimensional Koopman approximation and ensures the learned dictionary captures essential spectral features without relying on precomputed spectra or post-processing.
In this work, we parameterize the dictionary functions $\mathbf{\Psi}(x;\theta)$ using a simple feedforward neural network, as shown on the right side of Figure 1, where $\theta$ denotes the network parameters. Unlike EDMD, which minimizes prediction error, our method explicitly targets spectral accuracy.
Remark 3.1.
Notice that the neural network architecture can be replaced by more advanced models or other optimizers depending on the system and data structure.
Figure 1: The optimizer on the right looks for the optimal dictionary by minimizing Eq. (3.3).
In ResDMD’s framework, the spectral residual in Eq. (3.3) measures how well a computed eigenpair from the dataset, in other words, ideally we want the total spectral residual $J=\sum_{i=1}^{N_{K}}\widehat{\textit{res}}(\lambda_{i},\phi_{i})^{2}$ approaches to zero, which implies that we can treat $J$ as a loss function. Moreover, minimizing $J$ is also equivalent to the following minimization problem:
$\min_{\tilde{K}}J=\min_{\tilde{K}}\frac{1}{m}\|(\Psi_{Y}-\Psi_{X}\tilde{K})V\|_{F}^{2},$ (3.5)